find each probability if you pick one marble,replace it, then pick a second marble out of a bag of six mables with one black, two white, and 3 striped. what is the p of (striped, stiped)
Answer by stanbon(48568)

find each probability if you pick one marble,replace it, then pick a second marble out of a bag of six mables with one black, two white, and 3 striped. what is the p of (striped, stiped)
-----------
If you select "with replacement", the selections are independent.
P(striped, striped) = P(striped)*P(striped) = 3/6*3/6 = 1/4
==============
Cheers,
Stan H.